### v2

# PrunedLoRA: Robust Gradient-Based structured pruning for Low-rank Adaptation in Fine-tuning

Abstract
Low-rank adaptation (LoRA) has become a widely used paradigm for parameter-efficient fine-tuning of large language models, yet its representational capacity often lags behind full fine-tuning. Within the context of LoRA, a key open question is how to obtain expressive low-rank adapters from over-parameterized spaces. We propose PrunedLoRA, a new framework that leverages structured pruning to obtain highly representative low-rank adapters from an over-parameterized initialization. Unlike prior approaches that impose a fixed low-rank budget, PrunedLoRA dynamically prunes less important components during fine-tuning and prevents their reactivation, enabling flexible and adaptive rank allocation. For structured pruning, by minimizing the pruning error for overall loss, we provide fine-grained pruning and recovery updates in a gradient-based pruning strategy with grounded interpretation. We provide the first theoretical analysis of the robustness of structured pruning and provably show that under the impact of weight perturbation, gradient-based pruning is more robust than activation-based pruning with respect to overall loss. Empirically, PrunedLoRA consistently outperforms LoRA and its variants across supervised fine-tuning tasks in mathematical reasoning and natural language understanding, and it also demonstrates advantages over existing structured pruning methods across diverse sparsity levels.
1 Introduction
Low-rank adaptation (LoRA) [26] and its variant [92, 43, 23] have emerged as a prominent class of parameter-efficient fine-tuning (PEFT) methods for large- scale foundation models [59, 47, 96]. By injecting trainable low-rank matrices into the pre-trained model, LoRA enables efficient fine-tuning with minimal training overhead and no additional inference latency. Despite its efficiency, LoRA often lags behind full fine-tuning (FFT) in practical performance. Existing attempts to bridge this gap fall into two categories. The first line of work strictly follows LoRA’s memory constraint, so exploring over the full parameter space is inadmissible [23, 85, 29, 3]. Learning within the low-rank space is always difficult to utilize the powerful representation of FFT [93, 20]. The second line of work enables full-parameter learning [94, 24, 42] through projection techniques to compress and decompress gradients and weights. While these over-parameterized methods improve performance , they ultimately output fine-tuned full models rather than preserving a shared base model with lightweight, task-specific low-rank adapters. As a result, for the inference period, these approaches with full-parameter learning are less efficient, since each task requires storing a full model. In contrast, if we obtain low-rank adapters for different tasks, inference time and memory cost can be significantly reduced [81, 41, 14]. Therefore, the key question remains open: allowing for the cost of full-parameter learning [94], how can we find highly representative low-rank adapters from an over-parameterized setting to retain inference efficiency?
Empirically, we observe that increasing the rank of LoRA improves performance, in some cases approaching that of FFT (see Fig. 1 in Subsection 3.1), a trend also reported in prior work [72, 26]. This suggests that LoRA with a larger rank has sufficient representational capacity. Motivated by this observation, we consider initializing LoRA with a larger rank to ensure sufficient representational capacity, and then reducing the size of the model during fine-tuning to obtain a lightweight low-rank adapter. This strategy preserves the expressive power of an over-parameterized initialization while maintaining inference efficiency.
To realize this idea, we next turn to structured pruning [36, 22, 9, 97], a principled approach for reducing the model size by removing entire sub-components, such as rows or columns, from the model’s weight matrices. Two main categories of structured pruning have been widely studied: gradient-based methods [52, 83, 48] and activation-based methods [16, 32, 95]. Empirical evidence (e.g., [54]) suggests that gradient-based approaches focus more on global information and would be more stable for overall loss under weight perturbations. However, from a theoretical perspective, a clear comparison between these two classes of methods, particularly regarding how weight perturbations affect the overall loss, remains largely unexplored. To further mitigate the influence of pruning, [16, 32, 60] proposes updating weights after pruning, inspired by Optimal Brain Surgeon [21]. While these approaches investigate how to scale second-order methods to deep neural networks, they, as the original work [21], leave open a deeper understanding of the pruning metric, known as “saliency" term in Optimal Brain Surgeon.
In this work, aiming to obtain a low-rank adaptation at the end of post-training, we propose PrunedLoRA, enabling full-parameter learning while dynamically pruning the initial weights from an over-parameterized space. Unlike existing methods focusing on a fixed low-rank budget, PrunedLoRA enjoys the freedom of learning from over-parameterized spaces while converging to lightweight low-rank adapters for inference efficiency. For the theoretical analysis of structured pruning, we consider a toy model of self-attention [69] and provably show that gradient-based pruning is more robust to weight perturbations in terms of overall loss than activation-based pruning approaches. We further show that this intuition extends to broader contexts. In addition, we provide a fine-grained analysis of pruning selection and weight update for weight matrices in a second-order gradient-based pruning strategy, which deepens the understanding of the pruning metric (the “saliency" term in Eq. 5 of [21]) in the class of second-order pruning methods.
We summarize our contribution as follows:
  • We propose PrunedLoRA, a new framework that identifies highly representative low-rank adapters by structured pruning from an over-parameterized initialization with more representation capacity while retaining inference efficiency. Unlike prior approaches with a fixed low-rank budget, PrunedLoRA only enforces the low-rank constraint at the end of fine-tuning, enabling flexible and adaptive rank allocation during fine-tuning.
  • We establish the first theoretical analysis of the robustness of two major structured pruning approaches for large language models. Using a toy self-attention model, we prove that gradient-based pruning is more robust to weight perturbations in terms of overall loss than activation-based pruning, and we also show that this intuition extends to broader settings.
  • We conduct extensive experiments across supervised fine-tuning tasks spanning mathematical reasoning and natural language understanding, showing that PrunedLoRA can further narrow the gap between LoRA and FFT. Across different sparsity levels from $50\%$ to $93\%$ and across various pruning tasks (including both dynamic and one-shot pruning), our method consistently outperforms existing structured pruning methods.
2 Related Work
Low-rank adaptation (LoRA) has been extensively investigated in foundation models [4, 73, 40, 2, 96], with numerous variants and enhancements proposed [49, 23, 71]. [26] assumes that the fine-tuning update can be effectively captured in a low-rank subspace. Specifically, for a pre-trained model with weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, LoRA reparameterizes the weight update $\Delta{\bm{W}}$ via a low-rank decomposition as ${\bm{W}}_{0}+\Delta{\bm{W}}={\bm{W}}_{0}+s{\bm{B}}{\bm{A}}$, where ${\bm{B}}\in\mathbb{R}^{m\times r}$, ${\bm{A}}\in\mathbb{R}^{r\times n}$ and $s=\frac{\alpha}{r}$ is a scaling factor. Here, $r\ll\textit{min}(m,n)$ is the rank of the update. AdaLoRA [92] dynamically allocates the parameter budget by assigning more capacity to task-critical modules, but remains constrained within a limited subspace and does not fully explore the parameter space as in full fine-tuning. LoRA-Prune [91] leverages gradients from LoRA modules rather than the entire model to prune the whole model, which differs from our goal and leads to substantial performance degradation. In contrast, we only prune the trainable parameters to produce representative low-rank adaptations at the end.
Compression of Large Language Model (LLM) has gained a lot of attention and has been widely applied for parameter efficiency and reducing the latency [34, 65]. To compress the language model, previous works can be divided into several categories: network pruning [31, 79, 44, 18], knowledge distillation [62, 63, 56], quantization [84, 1, 89] and other techniques, like early exit [78]. In this work, we focus on structurally network pruning [38] to remove the entire filter from the neural network, whose approaches can be mainly categorized into two lines: activation-based pruning and gradient-based pruning. For the activation-based pruning [10, 27], it explores structured pruning based on activation statistics of neuron/filter output. If we aim to prune the weight matrix $\bm{W}$}, many activation-based strategies [16, 32, 77, 74] focus on the following optimization problem
[…]
where $\mathcal{{\bm{C}}}$ is a certain sparse structure. Inspired by Optimal Brain Surgeon [21], finding the optimal $\widehat{{\bm{W}}}$ in (1) takes two steps: find the optimal pruning column first and update the unpruned column [66, 32, 37]. For gradient-based strategies, by allowing access to the gradient of the overall loss, to measure the importance of $i$-th column in $\bm{W}$}, [92, 83] estimate the change in loss $\mathcal{L}$ once pruning the $i$-th column:
$I_{{\bm{W}}_{i}}=|\Delta\mathcal{L}_{{\bm{W}}_{i}}|=|\mathcal{L}_{{\bm{W}}_{i}}-\mathcal{L}_{{{\bm{W}}_{i}}=0}|.$ (2)
Here, computing the important score can help to find the pruned column, but it keeps the unpruned weight unchanged, without compensating for the influence of pruning. Thus, for a weight matrix, how to minimize the influence of pruning in gradient-based methods is important.
3 Methods
3.1 Motivation
Figure 1: Performance of standard LoRA [26] on GSM8K [7] with different ranks compared to full fine-tuning. Note that the method of full fine-tuning does not involve the initial rank, and we draw a red line here solely for comparison.
Motivation 1: Higher rank results in better performance. As illustrated in Figure 1, employing higher ranks in LoRA consistently leads to improved empirical performance on both GSM8K (see Sec. 4 for details). Notably, as the rank increases, the performance gradually converges toward that of full fine-tuning. This observation motivates our approach: rather than fixing LoRA to a small rank at the outset, we initialize with a sufficiently large rank—providing a number of trainable parameters close to full fine-tuning—and then progressively prune it to a smaller rank. Such a strategy may preserve most of the performance gains in over-parameterized settings while ultimately producing a memory-efficient low-rank adaptation.
Motivation 2: $\bm{A}$} and $\bm{B}$} in LoRA control the low-rank spaces. For the sub matrices, ${\bm{A}}\in\mathbb{R}^{r\times n}$ and ${\bm{B}}\in\mathbb{R}^{m\times r}$, we observe that the columns of $\bm{B}$} correspond to the column space of the original update $\Delta{\bm{W}}$, while the rows of $\bm{A}$} represent the row space [88]. Therefore, they can capture the row-wise and column-wise correlation separately. As we will discuss in the next section, pruning on sub-modules instead of the full matrix reduces the computational cost and simplifies the second-order structured pruning significantly.
3.2 The Robustness of Gradient-based structured pruning
Activation-based v.s. Gradient-based structured pruning. Pruning induces perturbations to the weights across layers of large language models, which in turn modifies the overall loss and may lead to a deterioration of empirical performance [16, 80]. Within the context of structured pruning [45, 53, 11], activation-based solving Problem (1) and gradient-based pruning using important scores in (2) are two main lines of approaches to find the optimal pruned structure. Intuitively, gradient-based methods focus more on the global correlation [54], so they shall be more robust for the overall loss under the influence of weight perturbation. However, no theoretical analysis provably shows the insight. Here, we analyze the influence of different pruning strategies on the overall loss. We provide formal analysis and general discussion in Appendix 7.
Proposition 1 (Unofficial Statement)
Suppose that, under activation-based and gradient-based pruning strategies, each module in a single attention module satisfies a given perturbation error. The error in the loss function would be linear w.r.t. perturbation error under different pruning strategies, but the error of activation-based methods depends on the magnitude of each module.
Proposition 1 reveals that the activation-based methods introduce a higher infatuation for the overall loss. It is consistent with the insight that activation-based methods cannot indicate the influence of weight change for global correlation [8]. Within the context of gradient-based pruning strategies, we formulate our problem on pruning the columns of a full weight matrix first. It would be interpreted as pruning the columns of matrix $\bm{B}$} (or the rows of matrix $\bm{A}$}) alone.
Problem formulation Our approach starts from the idea of applying a structured compression layer-wise, in a way that allows the layers to preserve most of their output characteristics. This setup is popular in the post-training quantization and unstructured pruning literature [16, 66, 75], and can be implemented as follows. In the fine-tuning period, the gradient is non-trivial as it helps the fine-tuned model align with the down-task data. Therefore, our setup is different from the literature in gradient-based pruning [60, 31]. We consider the perturbation of a single weight matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$ in a large language model. The pruned matrix is denoted as ${\bm{W}}+{\bm{\delta}}$, where the perturbation ${\bm{\delta}}\in\mathbb{R}^{m\times n}$ corresponds to pruning the same weight indices across all rows, i.e., entire columns are removed. The update ${\bm{\delta}}\in\mathbb{R}^{m\times n}$ is subject to the constraint that
${\bm{\delta}}_{:,\mathcal{M}_{s}}=-{\bm{W}}_{:,\mathcal{M}_{s}}.$ (3)
Here, $\mathcal{M}_{s}$ denotes the pruning mask that specifies the pruned column indices with sparsity s. Expanding the overall loss of the pruned model with weight matrix ${\bm{W}}+{\bm{\delta}}$ around $\bm{W}$} yields
$\begin{split}\mathcal{L}({\bm{W}}+{\bm{\delta}})&\approx\mathcal{L}({\bm{W}})+\left\langle\nabla_{\bm{W}}\mathcal{L}({\bm{W}}),{\bm{\delta}}\right\rangle+\frac{1}{2}tr(\textit{vec}({\bm{\delta}})^{\top}\mathbf{H}\,\textit{vec}({\bm{\delta}})),\end{split}$ (4)
which corresponds to the matrix-form second-order Taylor expansion, where $\textit{vec}({\bm{\delta}})$ denotes the vectorization of the perturbation matrix. Noticeably, the Hessian matrix is ${\bm{H}}\in\mathbb{R}^{mn\times mn}$, so the memory cost and the computational cost are extremely huge. To address the challenge, many existing methods propose to impose structural assumptions for the Hessian matrix $\bm{H}$}, such as diagonal or block-diagonal approximation [90, 21] and empirical Fisher [5, 60]. With the goal of selecting columns in (3), it is critical to preserve the correlation among the columns of the weight matrix. Thus, with the standard assumption of row independence in [31, 16], as a common technique for approximating the Hessian using gradients, we can approximate (4) by
$\begin{split}\mathcal{L}({\bm{W}}+{\bm{\delta}})&\approx\mathcal{L}({\bm{W}})+\left\langle\nabla_{{\bm{W}}}\mathcal{L}({\bm{W}}),{\bm{\delta}}\right\rangle+\frac{1}{2}tr({\bm{\delta}}^{\top}\widehat{{\bm{H}}}{\bm{\delta}}),\end{split}$ (5)
where $\widehat{{\bm{H}}}=(\nabla_{{\bm{W}}}\mathcal{L}({\bm{W}}))^{T}\nabla_{{\bm{W}}}\mathcal{L}({\bm{W}})\in\mathbb{R}^{n\times n}$. Then, combining the pruned structure (3) with the analysis of perturbation in $\bm{W}$}, it yields the optimal pruning selection and weight update by solving the following problem:
$\begin{split}\mathcal{M}_{s},{\bm{\delta}}=\textit{argmin}_{\mathcal{M}_{s},{\bm{\delta}}}&\left\langle\nabla_{{\bm{W}}}\mathcal{L},{\bm{\delta}}\right\rangle+\frac{1}{2}tr({\bm{\delta}}\widehat{{\bm{H}}}\,{\bm{\delta}}^{\top})\\ \textit{s.t.}\quad&{\bm{\delta}}_{:,\mathcal{M}_{s}}=-{\bm{W}}_{:,\mathcal{M}_{s}}.\end{split}$ (6)
Here, for simplicity, we denote $\nabla_{{\bm{W}}}\mathcal{L}({\bm{W}})$ as $\nabla_{{\bm{W}}}\mathcal{L}$. The optimal solution of $\bm{\delta}$} in ( 6) is
$\begin{split}{\bm{\delta}}&=-\nabla_{{\bm{W}}}\mathcal{L}\>\widehat{{\bm{H}}}^{-1}-{\bm{W}}_{:,\mathcal{M}_{s}}\left((\widehat{{\bm{H}}}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\right)^{-1}(\widehat{{\bm{H}}}^{-1})_{\mathcal{M}_{s},:}\\ &\quad+(\nabla_{{\bm{W}}}\mathcal{L}\widehat{{\bm{H}}}^{-1})_{:,\mathcal{M}_{s}}\left((\widehat{{\bm{H}}}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\right)^{-1}(\widehat{{\bm{H}}}^{-1})_{\mathcal{M}_{s},:}.\end{split}$ (7)
Interpretation for Algorithm Design. Let us further analyze the update $\bm{\delta}$} in (7). The first term in $\bm{\delta}$} is a second-order Newton step. If there is no sparse masking, it would be the optimal update utilizing second-order momentum. As $P_{\mathcal{M}_{s}}{\bm{\delta}}$ will only leave the second term in (7), which is a projection correction to ensure the pruned weights remain zero. Interestingly, it is dependent on the current weight $\bm{W}$} and the mask $\mathcal{M}_{s}$ but independent of the gradient $\nabla_{{\bm{W}}}\mathcal{L}$. The third term in (7) provides a dual variable compensation that projects the unconstrained Newton step into the feasible region. Once we get the closed-form solution of $\bm{\delta}$} in (7), the pruning problem in (13) is
$min_{\mathcal{M}_{s}}\operatorname{tr}\left((\mathbf{W}-\nabla_{{\bm{W}}}\mathcal{L}\>\mathbf{H}^{-1})_{:,\mathcal{M}_{s}}\>\left((\mathbf{H}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\right)^{-1}\>(\mathbf{W}-\nabla_{{\bm{W}}}\mathcal{L}\>\mathbf{H}^{-1})_{:,\mathcal{M}_{s}}^{\top}\right).$ (8)
Here, the pruning problem in (8) is closely related to the “saliency" term in [21]. With the analysis of matrix weight, we provide an explicit interpretation for second-order pruning strategies: we select the pruning mask that removes the columns whose post-update (second-order Newton update) values are least important under the Hessian-weighted quadratic metric. Existing methods deriving from Optimal Brain Surgeon can not provide a grounded interpretation from the “saliency" term, as most of them focus on the specific problems such as (1) [16, 32] or only analyze the one-dimensional weight vectors [8, 60, 31]. Therefore, our analysis enriches the understanding of the class of second-order pruning methods.
We summarize our solution in Algorithm 2 and we present a schematic illustration of the workflow in the left of Figure 2. In each pruning step, the pruning indices are determined by the gradient and the estimated Hessian.
3.3 PrunedLoRA
Algorithm 1 PrunedLoRA: structured pruning for Low-rank Adapters from over-parameterized spaces. We prune LoRA matrices $(A,B)$ with column sparsity $s$ on $B$ (and corresponding row sparsity s on AA) given gradients $(\nabla_{A}\mathcal{L},\nabla_{B}\mathcal{L})$ and Hessian estimates $(\widehat{H}_{A},\widehat{H}_{B})$.
1:Step 1: Search pruning mask.
$\arg\min_{\mathcal{M}_{s}}\;\mathrm{tr}\!\left(\widetilde{B}_{:,\mathcal{M}_{s}}\big((\widehat{H}_{B}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\big)^{-1}\widetilde{B}_{:,\mathcal{M}_{s}}^{\top}\right)+\mathrm{tr}\!\left(\widetilde{A}_{\mathcal{M}_{s},:}^{\top}\big((\widehat{H}_{A}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\big)^{-1}\widetilde{A}_{\mathcal{M}_{s},:}\right),$
where $\widetilde{A}=A-\widehat{H}_{A}^{-1}\nabla_{A}\mathcal{L},\;\widetilde{B}=B-\nabla_{B}\mathcal{L}\,\widehat{H}_{B}^{-1}$.
2:Step 2: Compute optimal updates.
3: Given $\mathcal{M}_{s}$, compute
$\delta_{B}=-\nabla_{B}\mathcal{L}\,\widehat{H}_{B}^{-1}-\widetilde{B}_{:,\mathcal{M}_{s}}\big((\widehat{H}_{B}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\big)^{-1}(\widehat{H}_{B}^{-1})_{\mathcal{M}_{s},:},$
$\delta_{A}=-\widehat{H}_{A}^{-1}\nabla_{A}\mathcal{L}-(\widehat{H}_{A}^{-1})_{:,\mathcal{M}_{s}}\big((\widehat{H}_{A}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}}\big)^{-1}\widetilde{A}_{\mathcal{M}_{s},:}.$
4: Set $A\leftarrow A+\delta_{A}$,   $B\leftarrow B+\delta_{B}$.
5:Step 3: Update LoRA adapters with standard optimizer in fine-tuning.
6:Step 4: Iterate or finalize.
7: If multi-round pruning is desired, repeat Steps 1–3 until the target rank is reached. Otherwise, output $(A,B)$.
In this part, we propose our structured pruning strategy, termed PrunedLoRA. Inspired by Motivation 1, we dynamically prune adapters $\bm{A}$} and $\bm{B}$} from high-parameter spaces.
Different from prior work such as AdaLoRA [92], which enforces an average rank budget and dynamically selects ranks from a small predefined set (e.g., $\{2,4,8\}$). It always restricts the rank of the updated weight in low-rank spaces. Besides, structurally pruning the columns and rows of a full weight matrix causes high computational overhead, as we highlight in Eq. (4). However, with Motivation 2, we can efficiently detect the row-wise and column-wise correlation by pruning the low-rank spaces of $\bm{A}$} and $\bm{B}$} together. With the goal of reducing the rank of the matrix, structured pruning of the decomposed sub-modules would be more efficient.
With the standard argument in Sec 3.2, the pruning problem for low-rank adaptation $\bm{A}$} and $\bm{B}$} is
Figure 2: Left: schematic of the dynamic pruning process, where the gradient and estimated Hessian will determine pruned columns and update as shown in Algorithm 2. Right: design of PrunedLoRA, where both adapter matrices $\bm{A}$} and $\bm{B}$} are jointly pruned under a masking scheme.
$\begin{split}\textit{argmin}_{\mathcal{M}_{s},{\bm{\delta}}_{{\bm{A}}},{\bm{\delta}}_{{\bm{B}}}}&\left\langle\nabla_{{\bm{A}}}\mathcal{L},{\bm{\delta}}_{{\bm{A}}}\right\rangle+\frac{1}{2}tr({\bm{\delta}}_{{\bm{A}}}^{\top}\,\widehat{{\bm{H}}}_{{\bm{B}}}\,{\bm{\delta}}_{{\bm{A}}})+\left\langle\nabla_{{\bm{B}}}\mathcal{L},{\bm{\delta}}_{{\bm{B}}}\right\rangle+\frac{1}{2}tr({\bm{\delta}}_{{\bm{B}}}\widehat{{\bm{H}}}_{{\bm{B}}}\,{\bm{\delta}}_{{\bm{B}}^{\top}})\\ \textit{s.t.}&\quad({\bm{\delta}}_{{\bm{B}}})_{:,\mathcal{M}_{s}}=-{{\bm{B}}}_{:,\mathcal{M}_{s}},\quad({\bm{\delta}}_{{\bm{A}}})_{\mathcal{M}_{s},:}=-{{\bm{A}}}_{\mathcal{M}_{s},:}.\end{split}$ (9)
Here, the mask $\mathcal{M}_{s}$ simultaneously controls the column sparsity of $\bm{B}$} and the row sparsity of $\bm{A}$}. Consequently, the Hessian estimates $\widehat{{\bm{H}}}_{\bm{A}}$ and $\widehat{{\bm{H}}}_{\bm{B}}$ are computed with different purposes: to capture the column-wise correlations of $\bm{B}$} and the row-wise correlations of $\bm{A}$}, respectively. Following the standard derivation in Sec 3.2, our pruning strategy for reducing high-rank matrices $\bm{A}$} and $\bm{B}$} to a low-rank adaptation begins by determining the optimal pruning mask via
$\begin{split}argmin_{\mathcal{M}_{s}}\operatorname{tr}\left(\widetilde{{\bm{B}}}_{:,\mathcal{M}_{s}}((\widehat{{\bm{H}}}_{{\bm{B}}})^{-1}_{\mathcal{M}_{s},\mathcal{M}_{s}})^{-1}\widetilde{{\bm{B}}}_{:,\mathcal{M}_{s}}^{T}\right)+\operatorname{tr}\left(\widetilde{{\bm{A}}}^{T}_{\mathcal{M}_{s},:}((\widehat{{\bm{H}}}_{{\bm{A}}})^{-1}_{\mathcal{M}_{s},\mathcal{M}_{s}})^{-1}\widetilde{{\bm{A}}}_{\mathcal{M}_{s},:}\right),\end{split}$ (10)
where $\widetilde{{\bm{A}}}={\bm{A}}-(\widehat{{\bm{H}}}_{{\bm{A}}})^{-1}\nabla_{{\bm{A}}}\mathcal{L},\>\widetilde{{\bm{B}}}={\bm{B}}-\nabla_{{\bm{B}}}\mathcal{L}\>(\widehat{{\bm{H}}}_{{\bm{B}}})^{-1}$. After selecting the pruning indices, we update $\bm{A}$} and $\bm{B}$} as (11) to minimize the perturbation error in the loss.
$\begin{split}{\bm{\delta}}_{{\bm{B}}}&=-\nabla_{{\bm{B}}}\mathcal{L}\>\widehat{{\bm{H}}}_{{\bm{B}}}^{-1}-\widetilde{{\bm{B}}}_{:,\mathcal{M}_{s}}((\widehat{{\bm{H}}}_{{\bm{B}}}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}})^{-1}(\widehat{{\bm{H}}}_{{\bm{B}}}^{-1})_{\mathcal{M}_{s},:}\\ {\bm{\delta}}_{{\bm{A}}}&=-\widehat{{\bm{H}}}_{{\bm{A}}}^{-1}\>\nabla_{{\bm{A}}}\mathcal{L}-(\widehat{{\bm{H}}}_{{\bm{A}}}^{-1})_{:,\mathcal{M}_{s}}((\widehat{{\bm{H}}}_{{\bm{A}}}^{-1})_{\mathcal{M}_{s},\mathcal{M}_{s}})^{-1}\widetilde{{\bm{A}}}_{\mathcal{M}_{s},:}.\end{split}$ (11)
Complexity. For PrunedLoRA, the pruning procedure begins with an initial rank smaller than $\min\{m,n\}$ and progressively reduces the rank until reaching the target level. Since pruning is performed only for a limited number of steps, the additional cost introduced by the pruning operations remains moderate. In particular, once the rank has been reduced to a value significantly smaller than $\min\{m,n\}$, the computational overhead of matrix inversion $\mathcal{O}(r^{3})$ becomes lower than that of matrix multiplication, i.e., $\mathcal{O}(\max\{m^{2}r,n^{2}r\})$. Consequently, our method maintains a computational cost comparable to that of existing low-rank adaptation approaches [85, 88].
4 Experiment
In this section, we present extensive experiments to evaluate the effectiveness of PrunedLoRA across various tasks and models. With different levels of pruning sparsity, we assess its capabilities on supervised fine-tuning tasks in mathematical reasoning using the Llama-3-8B model [17], and natural language understanding on a T5-based model covered in Sec 4.1. Then we conduct ablation studies for the hyperparameters, pruning schedules, and more pruning baselines in Sec 4.2 and Appendix 8.4. In addition to conducting structured pruning to obtain low-rank adaptation in fine-tuning, one-shot pruning for compressing a pretrained model is crucial in the pre-LLM era [61] as well, but most of the work [19, 61, 16] is activation-based methods without awareness of the influence of weight perturbation on the overall loss function. We provide a simple gradient-based method as well in Appendix 9 without weight update. It supports the effectiveness of gradient-based pruning strategies for eliminating the impact of weight perturbation.
Baselines. We compare PrunedLoRA with several representative fine-tuning paradigms to demonstrate its effectiveness. The first baseline is Full Fine-Tuning, where all parameters are updated. While this approach typically achieves the best performance, it is computationally expensive and offers no gains in inference efficiency. A widely adopted alternative is vanilla LoRA [26], which reparameterizes the updates through low-rank adapters $\bm{A}$} and $\bm{B}$}, initialized with Gaussian noise for $\bm{A}$} and zeros for $\bm{B}$}. We further consider two prominent LoRA variants that modify the low-rank structure: DoRA [43], which enhances representational capacity via learnable magnitude scaling, and AdaLoRA [92], which adaptively prunes and reallocates ranks based on singular value decomposition (SVD) to better capture parameter importance under a fixed budget. These variants constitute the most widely used structural extensions of LoRA. Other approaches, such as PiSSA [49] and rsLoRA [29], are largely orthogonal to pruning and could, in principle, be integrated into PrunedLoRA, which we leave as a promising complementary direction.
In addition to fine-tuning baselines, we also compare against existing structured pruning approaches for low-rank adaptation. Gradient-based pruning includes our method, which jointly optimizes parameter updates and pruning structure, as well as the widely used importance-score pruning strategy (Eq. 2) employed in LLM-Pruner [48]. Activation-based pruning determines the pruning structure based on input activation statistics (Eq. 1), as exemplified by ZipLM [32] and SparseGPT [16]. We further include comparisons with other classical pruning strategies [61, 19], along with one-shot pruning, which are reported in Appendix 8.4.
4.1 Experiments on Supervised Fine-tuning
Model and Datasets. To evaluate the scalability of PrunedLoRA, we fine-tune Llama-3-8B on mathematical reasoning. Besides, we fine-tune a T5-based model on a natural language understanding task. The experimental setup in this study follows closely the protocols established in prior LoRA research [72, 71].
Math: We train our model on a 100k subset of MetaMathQA [86], a dataset bootstrapped from other math instruction tuning datasets such as GSM8K [7] and math [25], with higher complexity and diversity. We select data bootstrapped from the GSM8K training set and apply filtering. The accuracy is reported on the GSM8K evaluation set.
Beyond the mathematical reasoning task, we further evaluate natural language understanding by fine-tuning a T5-base model [57] on a subset of the GLUE benchmark [70], including MNLI, SST-2, CoLA, QNLI, and MRPC. Model performance is assessed using accuracy as the evaluation metric.
[…]
Table 1: Performance comparison of fine-tuning and pruning baselines on GSM8K benchmark for Llama-3-8B-Base Model. Bold indicates the best result, underline represents the second-best one. ($\uparrow$:  higher values indicate better performance)
Method Before (%) After (%) Training Time
Full FT 100.00 100.00 4h 23min
LoRA 0.84 0.84 2h 28min
DoRA 0.89 0.89 2h 34min
AdaLoRA 0.84 0.84 2h 41min
PrunedLoRA (init r = 128) 1.68 0.84 2h 33min
PrunedLoRA (init r = 256) 3.36 0.84 2h 46min
PrunedLoRA (init r = 512) 6.71 0.84 3h 21min
Table 2: Comparison of trainable parameter ratios (before and after pruning) and training time across different fine-tuning methods.
Memory and Time Costs. In Table 2, we compare the percentage of trainable parameters (before and after pruning) and training time of our methods with FFT, LoRA, DoRA, and AdaLoRA on the math task and Llama-3-8B model. As the step number of structured pruning is quite small in the overall fine-tuning step, we have a comparable training time.
Results on Natural Language Generation. Table 1 shows that PrunedLoRA outperforms on both GSM8K. Compared with vanilla LoRA, which lags far behind full fine-tuning (64.4 vs. 73.3 on GSM8K), PrunedLoRA substantially closes the gap, and with init $r=512$ it even matches or surpasses full fine-tuning (73.38 on GSM8K). Relative to structured pruning baselines such as SparseGPT and LLM-Pruner, our method consistently yields higher accuracy, indicating greater robustness. We also find that larger initialization ranks lead to better outcomes, confirming our motivation that starting from higher-rank spaces provides richer expressiveness before pruning down to the final budget. We further provide a more detailed comparison across different initialization ranks for each pruning strategy (SparseGPT, LLM-Pruner, and PrunedLoRA) in Table 5 (See Appendix 5). These results confirm that PrunedLoRA consistently benefits from pruning higher-rank initializations.
Method MNLI   $\uparrow$ SST2   $\uparrow$ CoLA   $\uparrow$ QNLI   $\uparrow$ MRPC   $\uparrow$ Average   $\uparrow$
[…]
Table 3: GLUE benchmark results with different adaptation methods. Best results are in bold, second-best are underlined. ($\uparrow$:  higher values indicate better performance).
Results on Natural Language Understanding. In Table 3, we report the GLUE benchmark results for different adaptation methods. Full fine-tuning remains the best baseline overall, achieving the best average score of 87.91. Our proposed PrunedLoRA method narrows the gap between low-rank adaptation and fine-tuning by increasing the initial rank.
4.2 Experiments on Ablation Study
We conduct extensive ablation studies to better understand the design choices in PrunedLoRA. Detailed results are summarized in Appendix B.
Initialization Rank and Scaling Factor. We find that both the initialization rank and the scaling factor $\alpha$ critically affect the performance in Table 4. For a fixed rank, setting $\alpha$ proportional to the initialization rank yields the most stable convergence. For example, on GSM8K with rank 128, accuracy improves from 67.8 ($\alpha$ = $r/2$) to 69.2 ($\alpha$= $r$), while larger values ($\alpha$ = $2r$) provide little additional gain. Increasing the initialization rank further enhances results, with the accuracy rising to 72.1 at $r$ = 512 ($\alpha$= $r$). These results confirm the effectiveness of high-rank initialization combined with proportional scaling $\alpha$.
Pruning Schedule. We also vary the pruning interval ($K1$) and the number of columns pruned per step ($K2$) in Table 6. Gradual pruning with moderate intervals is consistently superior: pruning every 10 steps with $K2$ = 2 achieves the highest accuracy, while aggressive pruning ($K2$ = 4) slightly hurts performance. This suggests that maintaining stability during rank reduction is critical. Besides gradually pruning in post-training, we can also train LoRA with a high rank to converge and do one-shot structure pruning to obtain a low-rank adaptation (Appendix 8.4).
Target Rank. Beyond the default rank budget 64 in LoRA, we also examine more aggressive compression (e.g., pruning to target rank in $\{8,16\}$). As expected, extreme pruning leads to performance degradation, but PrunedLoRA remains competitive with or better than activation-based and simple gradient-based baselines at the same target rank (see Appendix 8.3). This highlights the robustness of structured pruning with the awareness of the overall under the cases of extreme compression.
5 Conclusion
In this work, we introduced PrunedLoRA, a gradient-based structured pruning framework for obtaining efficient low-rank adapters from over-parameterized spaces. By formulating pruning as an optimization problem that explicitly minimizes the loss induced by weight perturbations, our method provides a theoretically grounded strategy for structured adapter compression. Comprehensive experiments on mathematical reasoning and natural language understanding demonstrate that PrunedLoRA consistently narrows the gap to full fine-tuning while retaining inference efficiency. Furthermore, across diverse sparsity levels, it achieves superior performance over existing structured pruning baselines, underscoring both its robustness and practical effectiveness.